Error Analysis.The error in measuring the diameter of a large tree in inches is normal(-0.25,0.5).Find the probability that the error is between -1 and 1 inches.
Answer by stanbon(75887)

Error Analysis.The error in measuring the diameter of a large tree in inches is normal(-0.25,0.5).Find the probability that the error is between -1 and 1 inches.
------
z(-1) = (-1--0.25)/0.5 = -0.75/0.5 = -1.5
---
z(1) = (1--0.25)/0.5 = 1.25/0.5 = 2.5
----
P(-1< x <1) = P(-1.5< z < 2.5) = normalcdf(-1.5,2.5) = 0.9270
=========
Cheers,
Stan H.
=========